OpenStudy (anonymous):

Ms. Davis loves a good cup of coffee. If she puts on a pot of coffee using coneshaped coffee filter of radius 6 cm and depth 10 cm, and the water begins to drip out through the hole at the bottom at a constant rate of 1.5 cm^3 per second, determine how fast the water level is falling when the depth is 8 cm.

OpenStudy (anonymous):

What is the formula for the volume of a cone?

OpenStudy (anonymous):

V=1/3 pi r^2 d

OpenStudy (anonymous):

V=1/3 pi r^2 d

OpenStudy (anonymous):

correct?

OpenStudy (anonymous):

Probably. I don't remember. So you'll have to come up with another equation of r in terms of d. Probably something from similar triangles.

OpenStudy (anonymous):

...

OpenStudy (anonymous):

Ok well you know that when d =10, r = 6, And the ratio of the two will be constant throughout the cone. \[\frac{d}{r} = \frac{10}{6} \implies r = \frac{10}{6d}\]

OpenStudy (anonymous):

So now rewrite the volume of the cone formula just in terms of d.

OpenStudy (anonymous):

V= 1/3 pi (10/6d)^2*d

OpenStudy (anonymous):

And simplify.

OpenStudy (anonymous):

\[V = \frac{100\pi}{36d}\]

OpenStudy (anonymous):

Now what?

OpenStudy (anonymous):

Now solve for d in terms of V.

OpenStudy (anonymous):

so i plug in d=10 into my new Volume?

OpenStudy (anonymous):

No, because we want to know how fast d is changing at d=8

OpenStudy (anonymous):

oh yeah... so I got 25/72 pi

OpenStudy (anonymous):

where do I plug my changing rate?

OpenStudy (anonymous):

Oh wait. I was wrong. We just need to take the derivative of our new volume equation with respect to time.

OpenStudy (anonymous):

\[\frac{d}{dt}V = \frac{d}{dt}[\frac{100\pi}{36d}]\]

OpenStudy (anonymous):

Don't forget the chain rule!

OpenStudy (anonymous):

How do I take the derivative of that?

OpenStudy (anonymous):

\[\frac{d}{dt}V = \frac{100\pi}{36} * \frac{d}{dt}[\frac{1}{d}]\] \[ = \frac{100\pi}{36} * (ln\ d)*d'\]

OpenStudy (anonymous):

And we know that the Volume is changing at a rate of 1.5 per second. So: \[1.5 = \frac{100\pi}{36}(ln\ d) * d'\] Solve for \(d'\) and plug in 8 for d.

OpenStudy (anonymous):

why did you multiplyby 1/d?

OpenStudy (anonymous):

Because you had \[\frac{100\pi}{36d} = \frac{100\pi}{36} * \frac{1}{d}\]

OpenStudy (anonymous):

I've gotta go have dinner. I'll bbl

OpenStudy (anonymous):

okay.

OpenStudy (anonymous):

are you back?

OpenStudy (anonymous):

Did you figure it out?